1. The following differences enter into the reconciliation of financial income and taxable income of Abbott Company for the year ended December 31, 2017, its first year of operations. The enacted income tax rate is 30% for all years.

Pretax accounting income $800,000

Excess tax depreciation 480,000

Litigation accrual 70,000

Unearned rent revenue deferred on the books but appropriately recognized in taxable income

60,000 Interest income from New York municipal bonds 20,000

1. Excess tax depreciation will reverse equally over a four-year period, 2018-2021.

2. It is estimated that the litigation liability will be paid in 2021.

3. Rent revenue will be recognized during the last year of the lease, 2021.

4. Interest revenue from the New York bonds is expected to be $20,000 each year until their maturity at the end of 2021.

(a) Compute taxable income for 2017.

 (b) Prepare a schedule of future taxable and (deductible) amounts for the transactions listed above and show how they lead to the ending balance in the DTA/DTL accounts for 2017. Future Taxable (deductible amounts) 2018 2019 2020 2021 Total

(c) Record the tax journal entry for 2017

(d) Report the tax portion of the income statement for 2017

(e) Suppose Abbott finds out at the beginning of January 2018 prior to releasing the 2017 financial statements that its major customer filed for bankruptcy. Based upon this information, there is a 60% chance Abbott will sustain tax losses into the future such that they will not be able to use any of the DTA. Would you record an entry based upon this new information, if so what would it be?

(f) Suppose Congress enacts a new tax rate in December 2017, effective beginning in 2020, the new tax rate will be 20%. Prepare a modified schedule of future taxable and (deductible) amounts for the transactions listed above and show how they lead to the ending balance in the DTA/DTL accounts for 2017. Show corresponding journal entry. 2018 2019 2020 2021 Total
Answer #1

a) Computation of taxable income for 2017

.

Particulars Amount in $
Pretax accounting income       8,00,000
Add: Disallowed items
Litigation accrual          70,000
Add : Rent income          60,000
Taxable income       9,30,000

Note: Assuming interest income recognised in the pre-tax accounting period correctly

b)

Future taxable income or decuctible expenses
Amount in $
Year 2018 2019 2020 2021 Total
Deductible expenses - Litigation accrual                    -                       -                       -            -70,000          -70,000
Deductible income - Rent income                    -                       -                       -            -60,000          -60,000
Taxable income - Depreciation       1,20,000        1,20,000        1,20,000        1,20,000        4,80,000
Total       1,20,000        1,20,000        1,20,000          -10,000        3,50,000
DTA/DTL as on 2017
DTL
Depreciation       4,80,000
DTA
Unearned rental income recognised in tax          60,000
Litigation accrual          70,000
Net Deferred liability balance       3,50,000
Deferred Tax @ 30%       1,05,000

c) Journal entry for 2017

Tax expenses Dr 2,79,000 Income statement
To Provision for tax Cr 2,79,000 Balance Sheet
Deferred tax expenses Dr 1,05,000 Income statement
To Deferred tax liability Cr 1,05,000 Balance Sheet

d) Disclosure of tax portion of the income statement 2017

Tax expenses

Current tax - $ 2,79,000

Deferred tax - $ 1,05,000

e) Yes, Abott will record following entry in 2017

Bad debt written off ........Dr (Profit and loss)

To Account receivable ... Cr (Balance sheet)

f)

Modified schedule of future taxable and (deductible) amounts
Year 2018 2019 2020 2021 Total
Deferred asset balance
Deductible expenses - Litigation accrual                    -                       -                       -          -70,000          -70,000
Deductible income - Rent income                    -                       -                       -          -60,000          -60,000
Sub total (A)                    -                       -                       -       -1,30,000       -1,30,000
Deferred Tax Asset @30% (C=A*30%)                    -                       -                       -                     -                       -  
Deferred Tax Asset @20% (D=A*20%)                    -                       -                       -          -26,000          -26,000
Total deferred tax balance as on 2017 (H=C+D)                    -                       -                       -          -26,000          -26,000
Deferred liability balance
Taxable income - Depreciation       1,20,000        1,20,000        1,20,000      1,20,000        4,80,000
Subtotal (B)       1,20,000        1,20,000        1,20,000      1,20,000        4,80,000
Deferred Tax Asset @30% (E=B*30%)          36,000            36,000                     -                     -              72,000
Deferred Tax Asset @20% (F=B*20%)                    -                       -              24,000          24,000            48,000
Total deferred tax balance as on 2017 (G=E+F)          36,000            36,000            24,000          24,000        1,20,000
Deferred tax liability (G-H)          36,000            36,000            24,000          50,000        1,46,000